# Linguistic and Embedding-Based Profiling of Texts generated by Humans and Large Language Models

### v2

@@ Appendix D Linguistic Features Visualization
In this section, Figure 11 and Figure 12 show respectively the mean and standard deviation of the linguistic features values across domain per each model in the dataset. The following linguistic features are considered: Text Length, Sentence Length, Morphological Complexity Index (MCI) for nouns, Morphological Complexity Index (MCI) for verbs, dependency depth, dependency length, word prevalence, type-token ratio, semantic similarity and emotionality. We present the results of our analysis of our feature collection in each domain. Notably, we appreciate a difference in variability between domains, especially when dealing with more “creative” domains such as Poetry and Book (See Figures 11 and 12). Especially, humans show greater differences in average text length within these domains, reflecting humans’ tendency to adhere to genre-specific constraints in more rigid domains such as abstracts.
D.1 Poetry Domain
In our set of linguistic features, poetry domain shows humans scoring highest in average text length and emotionality, while maintaining fairly simpler syntactic structures than most models. We can also notice how humans tend to score high in morphological complexity for verbs and nouns.
D.2 Books Domain
In our set of linguistic features, Books domain shows humans score highest in average text length. However, Humans score lower on emotionality then some models, and they employ more complex syntactic structures than in Poetry. Nevertheless, we notice how humans tend to score high in morphological complexity for verbs and nouns in this domain as well.
D.3 Abstracts Domain
In our set of linguistic features, abstracts domain shows clear different patterns in comparison with Poetry and Books domain. Humans score lower on average text length and emotionality then with the other two domains. We can notice how recent models (e.g., GPT4) behave extremely more consistently and more similar to humans than their older models (e.g., GPT2) in text length, sentence length, syntactic depth and length, and also in semantic similarity, where older models tend to be way more varied in term of content than newer.
D.4 Recipes Domain
In our set of linguistic features, recipes domain shows humans scoring lowest in dependency length, underlying the tendency of employing simpler syntactic structures than models. This tendency involves semantic similarity as well, where humans appear extremely more consistent in the semantic content than previous discussed domains.
D.5 Reddit Domain
In our set of linguistic features, Reddit domain shows humans scoring lower on morphological complexity than previous discussed domains. This could point to differences in language register (e.g., less formal vs more formal) between a user-based domain with no genre-specific constraints and more strict domains (Biber and Conrad, 2019). Again, newer models perform more similar to humans than their older version, especially in semantic similarity (e.g., GPT).
D.6 Reviews Domain
In our set of linguistic features, Reviews domain shows humans employ a slightly higher use of frequently known words than in the other discussed domains, as we can see in Figure 11.
D.7 Wikipedia Domain
In our set of linguistics features, in Wikipedia domain humans score very low in average text length, as well in morphological complexity and emotionality, in comparison with the other discussed domains. Clearly, we can appreciate the effect of very rigid genre-specific constraints and how humans have to strictly abide to them.
D.8 News Domain
In our set of linguistic features, news domain show humans scoring high in average text length, morphological complexity, and fairly high in emotionality in comparison to most models. Differently from other domains, humans also score the lowest in word prevalence, pointing to the use of less common words, and lower in semantic similarity, underlying a more varied semantic content.
@@ D.8 News Domain
Figure 11: Multi-panel heatmap displaying log-normalized feature values across different text domains and models.
Figure 12: Multi-panel heatmap displaying log-normalized feature standard deviation across different text domains and models.

### v3

@@ Appendix E Linguistic Features Visualization
Figure 14 shows the mean values of the linguistic features (standard deviation in Figure 1) across domains for each model in the dataset. The following linguistic features are considered: Text Length, Sentence Length, Morphological Complexity Index (MCI) for nouns, Morphological Complexity Index (MCI) for verbs, dependency depth, dependency length, word prevalence, type-token ratio, semantic similarity and emotionality. Notably, we observe a difference in variability between domains, especially when dealing with more “creative” domains such as Poetry and Books (See Figures 14 and 1). Especially, human-written texts show greater differences in average text length within these domains.
@@
Figure 14: Multi-panel heatmap displaying log-normalized feature values across different text domains and models
Poetry Domain
In the poetry domain, on aggregate, HWT score highest in average text length and emotionality, while maintaining fairly simpler syntactic structures than most models. We can also notice that human-written texts tend to score higher in morphological complexity for verbs and nouns.
Books Domain
In the books domain, human-written texts score highest in average text length. However, HWT score lower on emotionality than some models, and they employ more complex syntactic structures than in Poetry. Nevertheless, we notice that HWT tend to score high in morphological complexity for verbs and nouns in this domain as well.
Abstracts Domain
For the abstracts domain, the patterns are different in comparison with poetry and books domains. Human-written texts score lower on average text length and emotionality than the other two domains. We can notice how recent models (e.g., GPT4) behave extremely more consistently and more similar to HWT than their older models (e.g., GPT2) in text length, sentence length, syntactic depth and length, and also in semantic similarity.
Recipes Domain
In our set of linguistic features, the recipes domain shows that human-written texts score lowest in dependency length, underlying the tendency to employ simpler syntactic structures than MGT. This tendency involves semantic similarity as well, where HWT appear extremely more consistent in the semantic content than in previously discussed domains.
Reddit Domain
In the Reddit domain, HWT score lower on morphological complexity than in previously discussed domains. This could point to differences in language register (e.g., less formal vs more formal) between a user-based domain with no genre-specific constraints and more strict domains (Biber and Conrad, 2019). Again, newer models show more similar patterns to human-written texts than their older versions, especially in semantic similarity (e.g., GPT).
Reviews Domain
The reviews domain shows that human-written texts employ a slightly higher use of frequently known words than in the other discussed domains, as we can see in Figure 14.
Wikipedia Domain
In the Wikipedia domain, HWT score very low in average text length, as well as in morphological complexity and emotionality, in comparison with the other discussed domains. Clearly, we can observe the effect of very rigid genre-specific constraints and how humans have to strictly abide by them.
News Domain
The news domain shows that HWT score high in average text length, morphological complexity, and fairly high in emotionality in comparison to most models. Unlike in other domains, human-written texts also score lowest in word prevalence, indicating the use of less common words, and lower in semantic similarity, suggesting a more varied semantic content.
Appendix F Experimental Setup Details
All experiments were carried out with an NVIDIA A100 GPU (40GB memory) and standard CPU resources. We did not train or fine-tune any models; instead, we relied on pre-trained encoders purely for inference:
  • StyleDistance, based on RoBERTa-base with approximately 125M parameters.
  • paraphrase-MiniLM-L6-v2, a MiniLM encoder with approximately 22M parameters.
The calculation of linguistic features (e.g., syntactic depth, sentence length, morphological complexity, emotion intensity) was performed using spaCy (en_core_web_sm) and other lexicon-based methods on CPU. Only the transformer-based components (style embeddings and semantic similarity) were executed on the GPU, requiring about 3 GPU hours. The overall computational budget was therefore modest, with CPU-bound feature extraction accounting for most of the runtime.